1.Solve Cos y cos px + sin y sin px = p

2.Solve

(D^3+2D^2+D)y= e^2x + x^2 + sin2x




1
Expert's answer
2021-06-09T17:00:50-0400

1.


"\cos y\cos (px)+\sin y\sin (px)=p"

"\cos(px-y)=p"

"px-y=\cos^{-1}(p)"

"y=px-\cos^{-1}(p), p=\dfrac{dy}{dx}"


This is a Clairaut differential equation.

Differentiating with respect to "x," both sides 


"p=p+x\dfrac{dp}{dx}+\dfrac{1}{\sqrt{1-p^2}}\cdot\dfrac{dp}{dx}"

"(x+\dfrac{1}{\sqrt{1-p^2}})\dfrac{dp}{dx}=0"

Hence


"x=-\dfrac{1}{\sqrt{1-p^2}} \text{ or }\dfrac{dp}{dx}=0=>p=c"

"x=-\dfrac{1}{\sqrt{1-p^2}}=>1-p^2=\dfrac{1}{x^2 }, \ x<0"

"p^2=\dfrac{x^2-1}{x^2}"

"p=\pm\dfrac{\sqrt{x^2-1}}{x}"

Substitute

"y=\pm(\sqrt{x^2-1}-\cos^{-1}(\dfrac{\sqrt{x^2-1}}{x})), -1\leq x<0"




"p=\dfrac{dy}{dx}=\dfrac{x}{\sqrt{x^2-1}}-\dfrac{1}{x\sqrt{x^2-1}}=\dfrac{\sqrt{x^2-1}}{x}"


Putting "p=c" in the equation we have


"y=cx-\cos^{-1}(c)"


2.


"(D^3+2D^2+D)y=e^{2x}+x^2+\sin(2x)"




"y'=z, y''=z', y'''=z''"


"z''+2z'+z=e^{2x}+x^2+\sin(2x)"

Homogeneous Equation


"z''+2z'+z=0"

The characteristic (auxiliary) equation


"r^2+2r+1=0"

"r_1=r_2=-1"

"z_h=c_1xe^{-x}+c_2e^{-x}"

"z_p=Ae^{2x}+B\sin(2x)+C\cos(2x)"

"+(Mx^2+Nx+Q)"

"z_p'=2Ae^{2x}+2B\cos(2x)-2C\sin(2x)"

"+(2Mx+N)"

"z_p''=4Ae^{2x}-4B\sin(2x)-4C\cos(2x)+2M"

Then


"4Ae^{2x}-4B\sin(2x)-4C\cos(2x)+2M"

"+4Ae^{2x}+4B\cos(2x)-4C\sin(2x)+4Mx+2N"

"+Ae^{2x}+B\sin(2x)+C\cos(2x)+Mx^2+Nx+Q"

"=e^{2x}+x^2+\sin(2x)"

"A=\dfrac{1}{9}"

"-3B-4C=1"

"4B-3C=0"

"M=\dfrac{1}{2}"

"N=-2"

"Q=3"


"z_p=\dfrac{1}{9}e^{2x}-\dfrac{3}{25}\sin(2x)-\dfrac{4}{25}\cos(2x)"

"+\dfrac{1}{2}x^2-2x+3"

Therefore


"z(x)=c_1xe^{-x}+c_2e^{-x}+\dfrac{1}{9}e^{2x}"

"-\dfrac{3}{25}\sin(2x)-\dfrac{4}{25}\cos(2x)+\dfrac{1}{2}x^2-2x+3"

"\int xe^{-x}dx=-xe^{-x}+\int e^{-x}dx"

"=-xe^{-x}-e^{-x}+C"


"y(x)=c_5+c_3xe^{-x}+c_4e^{-x}+\dfrac{1}{18}e^{2x}"

"+\dfrac{3}{50}\cos(2x)-\dfrac{4}{50}\sin(2x)+\dfrac{1}{6}x^3-x^2+3x"